# RHINO: A large horn antenna for detecting the 21cm global signalRHINO: A large horn antenna for detecting the 21cm global signal

### v1

@@ Abstract
The sky-averaged brightness temperature of the 21cm line from neutral hydrogen provides a sensitive probe of the thermal state of the intergalactic medium, particularly before and during Cosmic Dawn and the Epoch of Reionisation. This ‘global signal’ is faint, on the order of tens to hundreds of millikelvin, and spectrally relatively smooth, making it exceedingly difficult to disentangle from foreground radio emission and instrumental artefacts. In this paper, we introduce RHINO, an experiment based around a large horn antenna operating from $60-85$ MHz. Horn antennas are highly characterisable and provide excellent shielding from their immediate environment, which are potentially decisive advantages when it comes to the beam measurement and modelling problems that are particularly challenging for this kind of experiment. The system also includes a novel continuous wave calibration source to control correlated gain fluctuations, allowing continuous monitoring of the overall gain level without needing to rapidly switch between the sky and a calibration source. Here, we describe the basic RHINO concept, including the antenna design, EM simulations, and receiver electronics. We use a basic simulation and analysis pipeline to study the impact of the limited bandwidth on recovery of physical 21cm global signal model parameters, and discuss a basic calibration scheme that incorporates the continuous wave signal. Finally, we report on the current state of a scaled-down prototype system under construction at Jodrell Bank Observatory.
@@ 1 Introduction
Neutral hydrogen is one of the few truly ubiquitous cosmic tracers. It exists from around the time of last scattering at redshift $z\approx 1090$ and persists as one of the very few luminous sources through the cosmic Dark Ages (Furlanetto et al., 2006; Pritchard & Loeb, 2008; Fialkov & Loeb, 2013; Mondal & Barkana, 2023). After that, it is ionised in a patchy, localised fashion around the first stars and galaxies during Cosmic Dawn ($z\lesssim 30$), eventually becoming completely ionised on large scales by the end of the Epoch of Reionisation ($z\lesssim 10$). After reionisation, it exists mostly in dense self-shielded clumps within galaxies, themselves tracers of the underlying cosmological matter distribution, up to the present day (Bharadwaj & Sethi, 2001; Battye et al., 2004; Pritchard & Loeb, 2008; Morales & Wyithe, 2010).
The redshifted 21cm radio line provides the distinctive observational signature of neutral hydrogen gas (Hi) across this entire timespan. It is variously seen in absorption and emission during early epochs, depending on the dominant coupling of the hydrogen gas to other processes, and is mostly seen in emission at late times post-reionisation (Pritchard & Loeb, 2008). The spatial variations of the 21cm signal, as a function of direction and frequency, trace the large-scale structure of the Universe, as well as the ionisation state of the intergalactic medium (IGM) at higher $z$ (Madau et al., 1997).
Mapping out the brightness temperature fluctuations generally requires large radio telescope arrays, with sensitivity on a range of angular and spectral scales (Pober et al., 2014; Bull et al., 2015; Murray et al., 2024). The fluctuating signal is quite different from the much brighter foregrounds, however, which are spectrally smooth except for modulating effects due to the chromatic response of the instruments. The foreground contamination can be isolated into a wedge region in Fourier space (Thyagarajan et al., 2015), leaving a cleaner ‘window’ in which the 21cm fluctuations should in principle dominate. This permits a ‘foreground avoidance’ approach, in which only the relatively uncontaminated region of Fourier space needs to be retained to make a detection of the 21cm signal.
The sky-averaged, or global, 21cm signal, is a coarser observable, consisting only of a mean brightness temperature as a function of frequency (Shaver et al., 1999). The spatial fluctuations of the 21cm emission are largely averaged out on solid angles that are a significant fraction of the full sky, leaving only the smooth frequency-dependence of the average signal. This has a series of distinctive peaks and troughs, corresponding to epochs when large-scale heating or cooling of the IGM occurred, or the hydrogen gas coupled to different physical processes that changed its spin temperature (Sethi, 2005; Pritchard & Loeb, 2010; Glover et al., 2014). The mean brightness temperature is given by
@@
The evolution of the ionised fraction, and background and spin temperatures, is generally expected to be quite smooth, with peak and trough features in the mean brightness temperature typically spanning widths of $\gtrsim 10$ MHz or so in most physical models (Cohen et al., 2020). This makes separation of the faint 21cm signal from foreground contamination much harder, as instrumental, terrestrial, and atmospheric artefacts can readily impose structure on similar frequency scales (Anstey et al., 2021; Shen et al., 2021; Tripathi et al., 2024; Wang et al., 2024; Pattison et al., 2024b). As a result, extremely accurate instrumental calibration and characterisation is needed to make global signal experiments feasible.
@@
This is the central challenge for global signal experiments, and a variety of approaches have been taken to try and reach the requisite levels of precision. Examples of contemporary global signal experiments include EDGES (Rogers & Bowman, 2012; Bowman et al., 2018a), SARAS (2 and 3) (Singh et al., 2018; Nambissan T. et al., 2021), REACH (de Lera Acedo et al., 2022), MIST (Monsalve et al., 2024), and PRIZM (Philip et al., 2019), along with several other past, planned, and proposed efforts (e.g. Rogers & Bowman, 2012; Peterson et al., 2014; Sokolowski et al., 2015). These differ in terms of their detailed implementation, but share some common features. First, all of them use compact antennas, either dipoles (e.g. blade dipoles for EDGES and MIST, shaped petal/hexagonal dipoles for PRIZM and REACH respectively), spherical and conical monopoles (SARAS-2 and SARAS-3 respectively), and conical log-spiral (REACH’s second antenna). Many of these have been optimised to provide a wideband and relatively frequency-independent response, i.e. the beam full-width at half-maximum (FWHM) is almost constant in frequency (e.g. Cumner et al., 2022). Most of them also require a ground plane, which can span a wide area, but is potentially also a source of systematic artefacts (Bradley et al., 2019; Spinelli et al., 2022). In terms of the receiver electronics, these experiments typically use some elaboration of a Dicke-switched design (Dicke, 1946), in which the system frequently switches between observing the antenna and observing one or more calibration loads. There is more variation in the number of internal calibration targets, and how reflections in the signal chain are measured (e.g. see Sun et al., 2024), as well as differences in data transport (e.g. whether RF over fibre links are used for example).
@@
These commonalities are both a manifestation of convergent design (a sign that certain design features work well and so tend to be adopted widely), and a potential risk. Reflections and standing waves have been identified as important possible sources of spurious signals that can mimic the 21cm signal, and it is difficult to fully rid dipoles with ground planes of these issues (although several mitigations have been put in place by the experiments). Sensitivity to the detailed properties of the ground plane, and even what lies beneath it, introduce difficult modelling challenges (Rogers et al., 2022; Spinelli et al., 2022). The broad beams of the compact antenna designs result in a general sensitivity to the surrounding environment, including beneath the antenna. While generally seen as a robust calibration strategy, Dicke switching can also have some drawbacks, e.g. in terms of needing rapid switching between different signals paths, which must be separately characterised.
@@
In this paper, we introduce the Remote HI eNvironment Observer (RHINO), an attempt to build a global signal experiment that has a high level of independence from the ‘convergent’ designs. The central feature of RHINO is that it uses a large horn antenna instead of a compact or resonant design. Horn antennas are well-known for being characterisable with great accuracy (Balanis, 1988; Bird et al., 2007), hence their use in a number of precision-calibrated applications over the years, e.g. for Cosmic Microwave Background experiments, plus lower frequency experiments like TRIS (Zannoni et al., 2008) and L-BASS (Zerafa, 2023). With some tuning of the design, they can also achieve low sidelobes, a wideband response, and low cross-polarisation. Since the horns themselves are essentially waveguides, rather than ‘active’ receiving surfaces, there is also an advantage in terms of shielding of the actual feed from the surrounding environment, including the ground, and there is no strict need to incorporate a ground plane.
@@
In another choice to promote design independence, we also include a continuous wave (CW) calibration source (Pollak et al., 2019). This injects a known signal of large amplitude but very narrow spectral width, and permits continuous monitoring during observations without needing to Dicke switch. This can be used to significantly suppress correlated gain fluctuations if they are sufficiently correlated in frequency.
@@
Latitude $53^{\circ}14^{\prime}10^{\prime\prime}$
Longitude $2^{\circ}18^{\prime}26^{\prime\prime}$
@@ 2.1 Basic principles
Horn antennas can be thought of as tapered waveguides that shape the incoming EM radiation in such a way that the impedance of the receiver is smoothly matched to the impedance of free space. The horn itself merely provides a spatially-varying conductive boundary, and is not itself an active receiving surface; a simple coaxial feed set far back into the waveguide is what actually absorbs the EM waves. As with waveguides, only certain discrete modes satisfy the boundary conditions within the horn, and one can analyse different horn geometries according to which modes are supported (Balanis, 1988).
The minimum size of a horn is first set by the cutoff frequency, $\nu_{c}$, of its terminal waveguide. For a rectangular waveguide to support a mode ${\rm TE}_{mn}$, the cutoff frequency is defined as (Bird et al., 2007)
@@
For a pyramidal (rectangular) horn of a given aperture diameter, it is possible to find a so-called ‘optimum horn’ with dimensions that maximise the antenna gain. The optimum H and E-plane aperture side lengths are related to the height of the horn flare section, $h$, by (Balanis, 1988)
@@
The optimum is only achieved at a particular wavelength. Setting the target frequency to 70 MHz ($\lambda\approx 4.3$m), for an H-plane side of length 7.2m, this implies $h=4$m (or 6m for the E-plane). The horn length also affects other properties of the antenna pattern. The curvature of the wavefront is changing as it propagates between aperture and waveguide, and if the flare section is not long enough, some curvature will remain across the aperture, leading to an appreciable phase error and (e.g.) increased sidelobes. Diffraction at the edge of the aperture also contributes to an increased backlobe and far sidelobes. At high frequencies, design features such as corrugations can be used to more gently taper the field and improve the antenna pattern shape, but these would be difficult and costly to manufacture into a very large horn. Other possibilities, such as a choke around the aperture to suppress diffraction, are likely to be more practical in the $\lesssim 100$ MHz range.
Figure 1: Dimensions of the pyramidal horn that was selected as the reference design. The flare section has an aperture of $7.3\times 6.0$m, an edge slant length of 5.099m, and an overall vertical height of 4.3m. The rectangular waveguide section has an aperture of $3.54\times 2.00$m and a height of 2.997m. The horn will be static, pointing at zenith, and will have a welded mesh conductive surface supported by a scaffold frame. The location of the feed is marked on the waveguide.
@@ 2.2 Design requirements
    First sidelobes below --20 dB (peak-normalised power).
@@
    Backlobe below --30 dB (peak-normalised power).
@@
The beam FWHM range was chosen to ensure a broad instantaneous field of view, so as to fairly sample the monopole on the sky while limiting the sensitivity to low elevations, where most difficult-to-model terrestrial emission is found. The return loss requirement defines the minimum desired bandwidth, from 65–80 MHz (21% fractional bandwidth). We target a broader band of 60–85 MHz (34% fractional bandwidth), but anticipate degraded sensitivity at the band edges due to the receiver bandpass (which includes an FM bandstop filter for example). This band corresponds to redshifts between 15.7 – 22.7, and was chosen to overlap substantially with the EDGES feature, without extending into the FM band. Extending the band to lower frequencies would result in a larger and increasingly difficult to manufacture structure. We study the impact of targeting a relatively small bandwidth on recovery of 21cm global signal model parameters in Sect. 5.
@@
Figure 2: Slices through the normalised antenna pattern, as a function of frequency. The horizontal lines show: (dotted) the $-3$ dB (half-maximum) point; (dashed) the target sidelobe level; and (dot-dash) the target backlobe level. The shaded vertical regions show the target beam FWHM at the bottom and top of the design band.
@@
Sidelobes are typically more chromatic than the mainlobe, and are harder to measure in the field. As such, sidelobes are a potentially important and uncertain source of spectral variation in the data. The sidelobe level target was chosen to try to keep them below 1%, so that beam modelling errors would be correspondingly less important (although not necessarily negligible). At higher frequencies, it has been possible to build horn antennas with very low sidelobe levels, approaching and even surpassing $-30$ dB (e.g. Lawrie & Peters, 1966; Yassin et al., 1993; Zannoni et al., 2008; Wuensche et al., 2022). This has been achieved through designs with longer flare sections (which helps reduce phase errors) and the use of internal corrugations to shape the field within the horn. While technically feasible at lower frequencies too, either of these options would add significant expense and practical difficulties in the manufacture and maintenance of the antenna. [^1] For concreteness, our goal is to restrict the materials and construction costs to substantially below US\$500k at 2024 prices. On the other hand, by limiting ourselves to simpler and relatively compact geometries, we can expect generally larger sidelobe levels. We note that the first sidelobe is the main focus here; as well as normally being the largest, for the large beamwidths we consider, it is also likely to be the only sidelobe above the horizon.
@@
Finally, the backlobe requirement is intended to minimise the sensitivity of the antenna to the properties of the ground. As discussed above, dependencies on the ground plane shape and size, and even the electrical properties of the underlying soil, can prove challenging to model for compact antennas such as dipoles (Spinelli et al., 2022; Pattison et al., 2024a). For horn antennas, the backlobe is mostly due to diffraction (e.g. see Lawrie & Peters, 1966; Clarricoats & Olver, 1984), and so can in principle be managed through modifications of the antenna optics, e.g. by adding chokes, corrugations, or screens at the aperture. More complex solutions such as loading the horn with a metamaterial have also been proposed (e.g. Qi et al., 2015; Polo-López et al., 2018).

### v2

# RHINO: A large horn antenna for detecting the 21cm global signalData Availability

@@ Abstract
The sky-averaged brightness temperature of the 21cm line from neutral hydrogen provides a sensitive probe of the thermal state of the intergalactic medium, particularly before and during Cosmic Dawn and the Epoch of Reionisation. This ‘global signal’ is faint, on the order of tens to hundreds of millikelvin, and spectrally relatively smooth, making it exceedingly difficult to disentangle from foreground radio emission and instrumental artefacts. In this paper, we introduce RHINO, an experiment based around a large horn antenna operating from 608560-85 MHz. Horn antennas are highly characterisable and provide excellent shielding from their immediate environment, which are potentially decisive advantages when it comes to the beam measurement and modelling problems that are particularly challenging for this kind of experiment. The system also includes a novel continuous wave calibration source to control correlated gain fluctuations, allowing continuous monitoring of the overall gain level without needing to rapidly switch between the sky and a calibration source. Here, we describe the basic RHINO concept, including the antenna design, EM simulations, and receiver electronics. We use a basic simulation and analysis pipeline to study the impact of the limited bandwidth on recovery of physical 21cm global signal model parameters, and discuss a basic calibration scheme that incorporates the continuous wave signal. Finally, we report on the current state of a scaled-down prototype system under construction at Jodrell Bank Observatory.
@@ 1 Introduction
Neutral hydrogen is one of the few truly ubiquitous cosmic tracers. It exists from around the time of last scattering at redshift $z\approx 1090$ and persists as one of the very few luminous sources through the cosmic Dark Ages (2006PhR...433..181F; 2008PhRvD..78j3511P; 2013JCAP...11..066F; 2023NatAs...7.1025M). After that, it is ionised in a patchy, localised fashion around the first stars and galaxies during Cosmic Dawn ($z\lesssim 30$), eventually becoming completely ionised on large scales by the end of the Epoch of Reionisation ($z\lesssim 10$). After reionisation, it exists mostly in dense self-shielded clumps within galaxies, themselves tracers of the underlying cosmological matter distribution, up to the present day (2001JApA...22..293B; Battye:2004re; 2008PhRvD..78j3511P; 2010ARA&A..48..127M).
The redshifted 21cm radio line provides the distinctive observational signature of neutral hydrogen gas (Hi) across this entire timespan. It is variously seen in absorption and emission during early epochs, depending on the dominant coupling of the hydrogen gas to other processes, and is mostly seen in emission at late times post-reionisation (2008PhRvD..78j3511P). The spatial variations of the 21cm signal, as a function of direction and frequency, trace the large-scale structure of the Universe, as well as the ionisation state of the intergalactic medium (IGM) at higher $z$ (1997ApJ...475..429M).
Mapping out the brightness temperature fluctuations generally requires large radio telescope arrays, with sensitivity on a range of angular and spectral scales (2014ApJ...782...66P; Bull:2014rha; 2024JOSS....9.6501M). The fluctuating signal is quite different from the much brighter foregrounds, however, which are spectrally smooth except for modulating effects due to the chromatic response of the instruments. The foreground contamination can be isolated into a wedge region in Fourier space (Thyagarajan:2015ewa), leaving a cleaner ‘window’ in which the 21cm fluctuations should in principle dominate. This permits a ‘foreground avoidance’ approach, in which only the relatively uncontaminated region of Fourier space needs to be retained to make a detection of the 21cm signal.
The sky-averaged, or global, 21cm signal, is a coarser observable, consisting only of a mean brightness temperature as a function of frequency (1999A&A...345..380S). The spatial fluctuations of the 21cm emission are largely averaged out on solid angles that are a significant fraction of the full sky, leaving only the smooth frequency-dependence of the average signal. This has a series of distinctive peaks and troughs, corresponding to epochs when large-scale heating or cooling of the IGM occurred, or the hydrogen gas coupled to different physical processes that changed its spin temperature (2005MNRAS.363..818S; 2010PhRvD..82b3006P; Glover2014). The mean brightness temperature is given by
@@
The evolution of the ionised fraction, and background and spin temperatures, is generally expected to be quite smooth, with peak and trough features in the mean brightness temperature typically spanning widths of $\gtrsim 10$ MHz or so in most physical models (2020MNRAS.495.4845C). This makes separation of the faint 21cm signal from foreground contamination much harder, as instrumental, terrestrial, and atmospheric artefacts can readily impose structure on similar frequency scales (2021MNRAS.506.2041A; 2021MNRAS.503..344S; 2024MNRAS.528.1945T; 2024ApJ...967...87W; 2024MNRAS.527.2413P). As a result, extremely accurate instrumental calibration and characterisation is needed to make global signal experiments feasible.
@@
This is the central challenge for global signal experiments, and a variety of approaches have been taken to try and reach the requisite levels of precision. Examples of contemporary global signal experiments include EDGES (2012RaSc...47.0K06R; 2018Natur.555...67B), SARAS (2 and 3) (2018ExA....45..269S; 2021arXiv210401756N), REACH (2022NatAs...6..984D), MIST (2024MNRAS.530.4125M), and PRIZM (2019JAI.....850004P), along with several other past, planned, and proposed efforts (e.g. 2012RaSc...47.0K06R; 2014arXiv1409.2774P; 2015PASA...32....4S). These differ in terms of their detailed implementation, but share some common features. First, all of them use compact antennas, either dipoles (e.g. blade dipoles for EDGES and MIST, shaped petal/hexagonal dipoles for PRIZM and REACH respectively), spherical and conical monopoles (SARAS-2 and SARAS-3 respectively), and conical log-spiral (REACH’s second antenna). Many of these have been optimised to provide a wideband and relatively frequency-independent response, i.e. the beam full-width at half-maximum (FWHM) is almost constant in frequency (e.g. 2022JAI....1150001C). Most of them also require a ground plane to suppress spectral artifacts caused by reflections from the ground (although see e.g. 2024ApJ...961...56M). These ground planes sometimes span a wide area or are raised off the ground to avoid edge effects and other complications, but are potentially also a source of residual systematic artefacts themselves (2019ApJ...874..153B; 2022MNRAS.515.1580S). In terms of the receiver electronics, these experiments typically use some elaboration of a Dicke-switched design (dicke1946measurement), in which the system frequently switches between observing the antenna and observing one or more calibration loads. There is more variation in the number of internal calibration targets, and how reflections in the signal chain are measured (e.g. see 2024Univ...10..236S), as well as differences in data transport (e.g. whether RF over fibre links are used for example).
@@
These commonalities are both a manifestation of convergent design (a sign that certain design features work well and so tend to be adopted widely), and a potential risk. Reflections and standing waves have been identified as important possible sources of spurious signals that can mimic the 21cm signal, and it is difficult to fully rid dipoles with ground planes of these issues (although several mitigations have been put in place by the experiments). Sensitivity to the detailed properties of the ground plane, and even what lies beneath it, introduce difficult modelling challenges (2022RaSc...5707558R; 2022MNRAS.515.1580S). The broad beams of the compact antenna designs result in a general sensitivity to the surrounding environment, including beneath the antenna. While generally seen as a robust calibration strategy, Dicke switching can also have some drawbacks, e.g. in terms of needing rapid switching between different signals paths, which must be separately characterised.
@@
In this paper, we introduce the Remote HI eNvironment Observer (RHINO), an attempt to build a global signal experiment that has a high level of independence from the ‘convergent’ designs. The central feature of RHINO is that it uses a large horn antenna instead of a compact or resonant design. Horn antennas are well-known for being characterisable with great accuracy (balanis1988horn; bird2007horn), hence their use in a number of precision-calibrated applications over the years, e.g. for Cosmic Microwave Background experiments, plus lower frequency experiments like TRIS (2008ApJ...688...12Z) and L-BASS (Zerafa). With some tuning of the design, they can also achieve low sidelobes, a wideband response, and low cross-polarisation. Since the horns themselves are essentially waveguides, rather than ‘active’ receiving surfaces, there is also an advantage in terms of shielding of the actual feed from the surrounding environment, including the ground, and there is no strict need to incorporate a ground plane.
@@
In another choice to promote design independence, we also include a continuous wave (CW) calibration source (2019MNRAS.489..548P). This injects a known signal of large amplitude but very narrow spectral width, and permits continuous monitoring during observations without needing to Dicke switch. This can be used to significantly suppress correlated gain fluctuations if they are sufficiently correlated in frequency.
@@
Latitude $53^{\circ}14^{\prime}10^{\prime\prime}$N
Longitude $2^{\circ}18^{\prime}26^{\prime\prime}$W
@@ 2 Horn antenna design
Figure 1: (Left): Dimensions of the pyramidal horn that was selected as the reference design. The flare section has an aperture of $7.3\times 6.0$m, an edge slant length of 5.099m, and an overall vertical height of 4.3m. The rectangular waveguide section has an aperture of $3.54\times 2.00$m and a height of 2.997m. The horn will be static, pointing at zenith, and will have a welded mesh conductive surface. The location of the feed is marked on the waveguide. (Right): Detailed design drawing showing the multi-panel construction, consisting of $2.4\times 1.2$ m mesh panels cut to size/shape and welded into steel angle frames. These are then bolted together to form a self-supporting structure.
@@ 2.1 Basic principles
Horn antennas can be thought of as tapered waveguides that shape the incoming EM radiation in such a way that the impedance of the receiver is smoothly matched to the impedance of free space. The horn itself merely provides a spatially-varying conductive boundary, and is not itself an active receiving surface; a simple coaxial feed set far back into the waveguide is what actually absorbs the EM waves. As with waveguides, only certain discrete modes satisfy the boundary conditions within the horn, and one can analyse different horn geometries according to which modes are supported (balanis1988horn).
The minimum size of a horn is first set by the cutoff frequency, $\nu_{c}$, of its terminal waveguide. For a rectangular waveguide to support a mode ${\rm TE}_{mn}$, the cutoff frequency is defined as (bird2007horn)
@@
For a pyramidal (rectangular) horn of a given aperture diameter, it is possible to find a so-called ‘optimum horn’ with dimensions that maximise the antenna gain. The optimum H and E-plane aperture side lengths are related to the height of the horn flare section, $h$, by (balanis1988horn)
@@
The optimum is only achieved at a particular wavelength. Setting the target frequency to 70 MHz ($\lambda\approx 4.3$m), for an H-plane side of length 7.2m, this implies h=4h=4m (or 6m for the E-plane). The horn length also affects other properties of the antenna pattern. The curvature of the wavefront is changing as it propagates between aperture and waveguide, and if the flare section is not long enough, some curvature will remain across the aperture, leading to an appreciable phase error and (e.g.) increased sidelobes. Diffraction at the edge of the aperture also contributes to an increased backlobe and far sidelobes. At high frequencies, design features such as corrugations can be used to more gently taper the field and improve the antenna pattern shape, but these would be difficult and costly to manufacture into a very large horn. Other possibilities, such as a choke around the aperture to suppress diffraction, are likely to be more practical in the $\lesssim 100$ MHz range.
@@ 2.2 Design requirements
    First sidelobes below -20 dB (peak-normalised power).
@@
    Backlobe below -30 dB (peak-normalised power).
@@
These requirements were set on a qualitative basis, rather than flowing in detail from the scientific objective of suppressing systematic contamination of the recovered 21cm signal. The latter would involve end-to-end simulations that are strongly contingent on a choice of foreground removal method for instance (although such design studies have been attempted for other experiments, e.g. 2022MNRAS.509.4679A). Nevertheless, it is useful to have a target for optimisation. We justify each requirement as follows. The beam FWHM range was chosen to ensure a broad instantaneous field of view, so as to fairly sample the monopole on the sky while limiting the sensitivity to low elevations, where most difficult-to-model terrestrial emission is found. The return loss requirement defines the minimum desired bandwidth, from 65–80 MHz (21% fractional bandwidth). We target a broader band of 60–85 MHz (34% fractional bandwidth), but anticipate degraded sensitivity at the band edges due to the receiver bandpass (which includes an FM bandstop filter for example). This band corresponds to redshifts between 15.7 – 22.7, and was chosen to overlap substantially with the EDGES feature, without extending into the FM band. Extending the band to lower frequencies would result in a larger and increasingly difficult to manufacture structure. We study the impact of targeting a relatively small bandwidth on recovery of 21cm global signal model parameters in Sect. 5.
@@
Figure 2: Slices through the normalised antenna pattern, as a function of frequency. The horizontal lines show: (dotted) the 3-3 dB (half-maximum) point; (dashed) the target sidelobe level; and (dot-dash) the target backlobe level. The shaded vertical regions show the target beam FWHM at the bottom and top of the design band.
@@
Sidelobes are typically more chromatic than the mainlobe, and are harder to measure in the field. As such, sidelobes are a potentially important and uncertain source of spectral variation in the data. The sidelobe level target was chosen to try to keep them below 1%, so that beam modelling errors would be correspondingly less important (although not necessarily negligible). At higher frequencies, it has been possible to build horn antennas with very low sidelobe levels, approaching and even surpassing 30-30 dB (e.g. 1966ITAP...14..605L; 1993ITAP...41..357Y; 2008ApJ...688...12Z; Wuensche:2021dcx). This has been achieved through designs with longer flare sections (which helps reduce phase errors) and the use of internal corrugations to shape the field within the horn. While technically feasible at lower frequencies too, either of these options would add significant expense and practical difficulties in the manufacture and maintenance of the antenna. [^1] For concreteness, our goal is to restrict the materials and construction costs to substantially below US\$100k at 2025 prices. On the other hand, by limiting ourselves to simpler and relatively compact geometries, we can expect generally larger sidelobe levels. We note that the first sidelobe is the main focus here; as well as normally being the largest, for the large beamwidths we consider, it is also likely to be the only sidelobe above the horizon.
@@
Finally, the backlobe requirement is intended to minimise the sensitivity of the antenna to the properties of the ground. As discussed above, dependencies on the ground plane shape and size, and even the electrical properties of the underlying soil, can prove challenging to model for compact antennas such as dipoles (2022MNRAS.515.1580S; 2025MNRAS.538.1301P). For horn antennas, the backlobe is mostly due to diffraction (e.g. see 1966ITAP...14..605L; clarricoats1984corrugated), and so can in principle be managed through modifications of the antenna optics, e.g. by adding chokes, corrugations, or screens at the aperture. More complex solutions such as loading the horn with a metamaterial have also been proposed (e.g. qi2015suppressing; polo2018compact).